A student knows 30 of 35 answers that should be learned for the exam. There two questions in a question card.what is the probability that student doesn't know all two answers?


1
Expert's answer
2021-01-07T18:10:31-0500

The total number of combination for preparing a question card with 2 questions out of 35 questions is given by 35C2


The total number of combinations for not knowing both the questions from a question card is given by 5C2 (since there are 5 questions (35-30=5) that a student does not know out of 35 questions)


Hence the probability that a student does not know both the questions in a question card is given by,


"\frac {5C2}{35C2} = \frac{\frac{5*4}{2*1}}{\frac{35*34}{2*1}}=\frac {5*4}{35*34}=0.0168"


The probability that student doesn't know all two answers = 0.0168